# Iterative Gaussian Approximation for Random Spreading Unsourced Random Access

### v1

@@ I-A Background and Related Work
The Gaussian multiple access channel (GMAC) is one of the most important application scenarios for URA [11], focusing on the design of basic MA concatenated codes. Its effectiveness has been validated under realistic channel conditions [12, Fig. 2b]. Several promising GMAC URA schemes have been proposed, including sparse codes [12][13][14][15], random spreading [16][17][18][19], segmented schemes [20][21, 22], among others. In particular, random spreading URA (RS-URA) shows strong anti-interference capabilities by leveraging the spectrum spreading structure.
@@ I-B Challenges and Contributions
Iterative channel decoding, a common technique in sparse codes [11], iterates extrinsic and intrinsic soft information between multi-user detection and the channel decoder to enhance decoding performance, as exemplified in [27, Eq. 5–Eq. 12]. This approach unlocks diversities within the channel code domain, providing valuable feedback for overall decoding, such as interference cancellation in multiple access systems. However, its application to RS-URA has been underexplored.
A recent attempt to leverage channel code domain diversities is described in [19, Section IV-B], where hard-verdict feedback information from the channel decoder updates the calculation of a posteriori constellation symbol probabilities, aiding multi-user detection in the corresponding algorithms. Nevertheless, this approach is constrained to a specific algorithm’s structure and serves as an additional procedure to improve estimation precision, lacking universality.
@@ III-B1 Active signature Detection
where $\mathbf{\Gamma}$ is a diagonal matrix with only elements corresponding to the active signatures are ones and other diagonal elemtns are zeros, and $\mathbf{X}$ is a row-sparse matrix with rows corresponding to the active users are $\mathbf{x}_{i}$’s and other rows are zero. One advantage of using the RS structure is that the received signal model can be expressed in the form of (4), which corresponds to a multiple measurement vector (MMV) model. This enables the application of robust compressive sensing (CS) algorithms to obtain favorable solutions. These include correlation-based greedy matching pursuit methods [17, 16, 18] or more advanced approximate message passing (AMP) algorithms [22, 21, 19], which are designed based on the signal’s statistical features. For a performance comparison between the correlation-based and AMP-based algorithms, refer to [19, Fig. 1].
@@ IV Numerical Results
    TABLE I: Signal amplitudes for different $K_{a}$ and group numbers $m$, derived from [27, Theorem 1]

### v2

@@ I-A Background and Related Work
The Gaussian multiple access channel (GMAC) is one of the most important application scenarios for URA [11, 12], focusing on the design of basic MA concatenated codes. Its effectiveness has been validated under realistic channel conditions [13, Fig. 2b]. Several promising GMAC URA schemes have been proposed, including sparse codes [13][14][15][16], random spreading [17][18][19][20], segmented schemes [21][22, 23], among others. In particular, random spreading URA (RS-URA) shows strong anti-interference capabilities by leveraging the spectrum spreading structure.
@@ I-B Challenges and Contributions
Iterative channel decoding, a common technique in sparse codes [11], iterates extrinsic and intrinsic soft information between multi-user detection and the channel decoder to enhance decoding performance, as exemplified in [12, Eq. 5–Eq. 12]. This approach unlocks diversities within the channel code domain, providing valuable feedback for overall decoding, such as interference cancellation in multiple access systems. However, its application to RS-URA has been underexplored.
A recent attempt to leverage channel code domain diversities is described in [20, Section IV-B], where hard-verdict feedback information from the channel decoder updates the calculation of a posteriori constellation symbol probabilities, aiding multi-user detection in the corresponding algorithms. Nevertheless, this approach is constrained to a specific algorithm’s structure and serves as an additional procedure to improve estimation precision, lacking universality.
@@ III-B1 Active signature Detection
where $\mathbf{\Gamma}$ is a diagonal matrix with only elements corresponding to the active signatures are ones and other diagonal elemtns are zeros, and $\mathbf{X}$ is a row-sparse matrix with rows corresponding to the active users are $\mathbf{x}_{i}$’s and other rows are zero. One advantage of using the RS structure is that the received signal model can be expressed in the form of (4), which corresponds to a multiple measurement vector (MMV) model. This enables the application of robust compressive sensing (CS) algorithms to obtain favorable solutions. These include correlation-based greedy matching pursuit methods [18, 17, 19] or more advanced approximate message passing (AMP) algorithms [23, 22, 20], which are designed based on the signal’s statistical features. For a performance comparison between the correlation-based and AMP-based algorithms, refer to [20, Fig. 1].
@@ IV Numerical Results
    TABLE I: Signal amplitudes for different $K_{a}$ and group numbers $m$, derived from [12, Theorem 1]